OpenStudy (anonymous):

The Empire State Building reaches one thousand two hundred fifty feet at the top floor. Keira has a model of this landmark building that is 8 inches tall. What is the scale between the model and the actual building?

OpenStudy (anonymous):

help me now

OpenStudy (goldphenoix):

"Scale- The ratio of the length in a drawing (or model) to the length of the real thing. Example: in the drawing anything with the size of "1" would have a size of "10" in the real world, so a measurement of 150mm on the drawing would be 1500mm on the real horse." So the height of the Empire State Building is 1,250 ft. The height of the model is 8 inches. You first want to convert 1,250 ft into inches. Since 12 inches is in 1 feet. Then 1,250 * 12 would represent the height of the Empire State Building in inches. So, 1,250 * 12 = ?

OpenStudy (anonymous):

15000

OpenStudy (goldphenoix):

Now they're asking you, "the scale between the model and the actual building" Since, the model is 8 inches, and the actual building in inches is 15,000 inches. You can put that into a ratio. 8:15,000 Or 8/15,000. The final step is to simplify the ratio. Do you know how to simplify?

OpenStudy (anonymous):

no

OpenStudy (goldphenoix):

Well, you want to find the greatest common factor of 8 and 15,000. The easiest way to do it is to find any common factor of 8 and 15,000. Do you think 2 is a common factor of 8 and 15,000?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Great. So divide 8 and 2. And 15,000 and 2. 8/2 = 4. 15,000 / 2 = 7,500. Now we have 4:7,500 Is it in simpliest form? Well, is 2 a common factor of 4 and 7,500?

OpenStudy (anonymous):

so whats the answer

OpenStudy (goldphenoix):

Well, answer my question first. Is 2 a common factor of 4, and 7,500?

OpenStudy (anonymous):

what do I do

OpenStudy (goldphenoix):

Divide 4 by 2 And 7,500 by 2. What does that give you?

OpenStudy (anonymous):

4 by 2 is 2

OpenStudy (goldphenoix):

Good. And 7500 /2 = ?

OpenStudy (anonymous):

3750

OpenStudy (anonymous):

3750

OpenStudy (goldphenoix):

Good. Now we have: 2:3750. Can this be simplify more? Well, 2 / 2 =? 3750 / 2 = ?

OpenStudy (anonymous):

1875

OpenStudy (goldphenoix):

Great. Now the scale looks like 1:1875. It can't be simplfy more because the only common factor of 1 is 1. So the answer is: 1: 1875 or 1/1875.

OpenStudy (anonymous):

thx

OpenStudy (goldphenoix):

No problem. :) I hope this help.